OpenStudy (anonymous):

Solve. Please show all of your work. If James received a 9% raise and is now making $102,000 per year, what was his salary before the raise? Round your answer to the nearest dollar.

Parth (parthkohli):

102000 = x + 9x/100 102000 = 109x/100 10200000 = 109x => 93978 dollars

OpenStudy (mathmagician):

109%=102000 100%=x so x=100*102000/109=93578$

Parth (parthkohli):

I meant 93578..(exact value is 93 577.9817)

OpenStudy (anonymous):

ok thanks